| ID | 
Date | 
Author | 
Group | 
Subject | 
Text | 
  |  
| 
  68  
 | 
03 Mar 2023, 23:25 | 
Giuseppe, Sergii | SCS | Late shift summary | achievements:
 
 
 
	change the optical laser 
to 2 um
 
	found focus of 2 um beam
 
	found | 
   |  | 
  67  
 | 
03 Mar 2023, 23:06 | 
Ulf Zastrau | HED | Shift summary | Beam delivery SASE 8,25 keV, 1300 µJ. 
Intensity stable,
 
but beam jittery in horizontal.  | 
   |  | 
  65  
 | 
02 Mar 2023, 23:13 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	9.3 keV, 202 pulses, X-ray 
intensity had slope over train, 30 % difference | 
    |  | 
  66  
 | 
02 Mar 2023, 22:58 | 
Giuseppe Mercurio and Sergii | SCS | late shift summary | achievements:
 
 
 
	knife edge scans of OL 
to find lens position corresponding to the | 
   |  | 
  64  
 | 
02 Mar 2023, 22:54 | 
Ulf Zastrau | HED | Stability Issue | HED would like to report 
that since 2 days the SASE2 beam is instable.
 
The attached movie shows, on the left | 
    |  | 
  63  
 | 
02 Mar 2023, 05:41 | 
Teguh, Luigi | SCS | Night shift summary 02.03.2023 | Achievements:
 
 
1. Performed knife-edge scan. The 
beam size is currently at 5 um x 120 um. | 
   |  | 
  62  
 | 
01 Mar 2023, 23:31 | 
Sergii, Martin | SCS | Late shift summary 01.03.2023 | Achievments:
 
 
- Align DPS to remove clipping 
of X-rays | 
   |  | 
  61  
 | 
01 Mar 2023, 23:25 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	9.3 keV, 1-352 pulses, 
0.8 mJ, stable delivery
 
 
 
Optical | 
   |  | 
  60  
 | 
01 Mar 2023, 16:07 | 
Alexander Pelka | HED | Shift summary | Day shift 28.02
 
 
Got the 
beam, aligned it up to the TCC, mirrors in | 
   |  | 
  59  
 | 
28 Feb 2023, 23:24 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	9.3 keV, 1-352 pulses, 
0.8 mJ, stable delivery
 
 
 
Optical | 
   |  | 
  58  
 | 
28 Feb 2023, 07:17 | 
Romain Letrun | SPB | Issue | In the public status display, it looks 
like FXE is taking the beam even though their 
shutter is closed. | 
    |  | 
  57  
 | 
27 Feb 2023, 02:10 | 
Felix Brausse | MID | Shift Summary Feb 26th | Moved out MONO, aligned SDL without MONO
 
 
Looked at both beams on DES
 
 
Struggled to get the same beam | 
   |  | 
  56  
 | 
26 Feb 2023, 01:37 | 
Felix Brausse | MID | Shift Summary Feb 25th | 
 
	Aligned both branches of the split-and-delay 
line.
 
	Brought the beam into the 
experimental hutch to image it on the x-ray | 
    |  | 
  55  
 | 
24 Feb 2023, 22:57 | 
Alexey Zozulya |   | Shift summary | - Measured the speckle patterns using EPIX 
detector to estimate the beamsize focused 
down by 1D lens (Row 13). | 
   |  | 
  54  
 | 
24 Feb 2023, 05:39 | 
Adam Round | SPB | Shift summary | Staff: AR, KK
 
 
X-ray delivery
 
 
 
	9.3keV
 
 
 
Optical | 
   |  | 
  53  
 | 
24 Feb 2023, 00:55 | 
Jan-Etienne Pudell |   | shift summary | We tested pulse on demand scans with 
scantool and confirmed that there is only 
one shot per scan point. This was used to | 
   |  | 
  52  
 | 
23 Feb 2023, 01:51 | 
Jan-Etienne Pudell |   | shift summary | After realignement of M1 and M2 we could 
solve the problem of the elogated beamprofile. 
Different lens setting were prealignd for | 
   |  | 
  51  
 | 
22 Feb 2023, 14:53 | 
Peter Zalden | FXE | Stability Issue | Instability observed today at 14:05, reported 
to PRC and BKR, including RC: We saw RMS 
jitter amplitude of up to 4 mm at FXE imager. | 
    |  | 
  50  
 | 
21 Feb 2023, 23:31 | 
Thomas Baumann |   | SASE1 changing photon energy changes the 2-color setup in SASE3 | We were set-up for 5.7 keV photon energy 
in SASE1. We made a test with BKR to change 
this energy to 9.3 keV. This had an effect | 
     |  | 
  49  
 | 
21 Feb 2023, 23:07 | 
Jan-Etienne Pudell |   | shift summary | - Tunnel + experimental hutch alignment 
(Mirrors, CRL1 and CRL2)
 
 
- Aligned main devices: Optics | 
   |   
 |